OpenStudy (anonymous):

if dy/dx= (3sinx)/(sec^2 x), then y=

OpenStudy (anonymous):

integrate

OpenStudy (anonymous):

@sourwing integrate the whole thing or just dy part? cuz my answer key says y= -cos^3+2, i just dont know how

OpenStudy (anonymous):

Rewrite this as \(3 \int {\sin{x}\cos^2{x}} \,\,dx\).

OpenStudy (anonymous):

Recall that the derivative of \(\cos{x}\) is \(-\sin{x}\) and use \(u\)-substitution.

OpenStudy (anonymous):

@Richard_Feynman then what happen to the sec^2 x?

OpenStudy (anonymous):

Recall that \(\sec{x} = \frac{1}{\cos{x}}\).

OpenStudy (anonymous):

So \(\frac{1}{\sec^2x} = \cos^2x\)